OpenStudy (anonymous):

In the equation y equals one-fourth x plus 8, what is y when x is twenty? twenty-eight thirteen twelve 7

OpenStudy (charlotte123):

y = 1/4x + 8

OpenStudy (anonymous):

thirteen

OpenStudy (anonymous):

https://www.google.com/search?client=safari&rls=en&q=y=1/4%2B8&ie=UTF-8&oe=UTF-8#q=y%3D1%2F4x%2B8&rls=en you can look up the equation on google if you don't have a lot of time

OpenStudy (charlotte123):

x = 20 1/4 * 20 + 8 =

OpenStudy (charlotte123):

1/4 * 20 = 5 y = 5 + 8

OpenStudy (anonymous):

y=1/4 x 20 + 8 Remove the parentheses from the numerator. y=(20)/(4)+8 Reduce the expression (20)/(4) by removing a factor of 4 from the numerator and denominator. y=5+8 Add 8 to 5 to get 13. y=13